# Hiding Local Manipulations on SAR Images: a Counter-Forensic Attack

### v1

@@ VI-C Realistic Attacks
VI-D Why not simpler counter-forensic attacks?
In this section, we perform further experiments to validate the need of the proposed speckle injection step composing our counter-forensic pipeline. Specifically, we compare the proposed global processing operation based on speckle injection with more standard counter-forensic attacks that are usually employed to hinder forensic traces from digital images. We investigate if speckle injection is actually the best processing operation to conceal the manipulation traces. Therefore, instead of injecting a phase-only speckle noise into the manipulated amplitude images (see Section IV-B1), we pass the images to multiple different global operations alternative to speckle. After the global editing, we apply the same pipeline shown in Fig. 6: we pass each processed manipulated image through the estimated SAR end-to-end system, then we match the histogram with that of the input manipulated image.
We consider nine editing operations, listed in Table VI. In particular, the “UpDownScale” and “DownUpScale” operations consist in the application of consecutive upscaling and downscaling (or viceversa) with one resizing factor and its opposite, such to match the initial image size at the end of the processing.
@@ VI-D Why not simpler counter-forensic attacks?
Counter-forensic attack evaluation. We report the achieved detection results on the attacked images in Table VII. For simplicity, we only consider the ASAE detector, as it demonstrated the most sensitive to counter-forensic attacks (see Table V). All the alternative global operations report extremely high AUC values, meaning for absence of counter-forensic effect. The proposed speckle injection process demonstrates the best option to hide manipulation artifacts.
TABLE VI: Global editing operations considered during the ablation study. $\mathcal{N}$, $\mathcal{L}$, $\mathcal{P}$ and $\mathcal{U}$ indicate the Gaussian, Laplace, Poisson and Uniform distributions, respectively.
Global editing operation Parameters
Gaussian Blur $\sigma=0.5$
UpDownScale Near Resize factors =$=\{1.05,1/1.05\}$
DownUpScale Near Resize factors =$=\{1/1.05,1.05\}$
UpDownScale Far Resize factors =$=\{1.5,1/1.5\}$
DownUpScale Far Resize factors =$=\{1/1.5,1.5\}$
Additive Gaussian Noise $\sim\mathcal{N}(0,0.0005\cdot(2^{16}-1))$
Additive Laplacian Noise $\sim\mathcal{L}(0,0.0005\cdot(2^{16}-1))$
Additive Poisson Noise $\sim\mathcal{P}(0.0005\cdot(2^{16}-1))$
Additive Uniform Noise $\sim\mathcal{U}[-50,50]$
TABLE VII: Average AUC achieved in presence of counter-forensic attack, for different global editing operations considered in the ablation study. In bold, the lowest achieved AUC per editing operation.
Editing operation
Global Editing Gaussian Blur Upscale Near Upscale Far Downscale Near Downscale Far Rotate Near Rotate Far
Additive Gaussian $0.893$ $0.912$ $0.932$ $0.785$ $0.653$ $0.879$ $0.863$
Additive Laplacian $0.849$ $0.865$ $0.886$ $0.755$ $0.655$ $0.835$ $0.823$
Additive Poisson $0.902$ $0.923$ $0.945$ $0.787$ $0.649$ $0.888$ $0.870$
Additive Uniform $0.908$ $0.928$ $0.947$ $0.794$ $0.653$ $0.894$ $0.877$
Speckle Injection $\mathbf{0.621}$ $\mathbf{0.636}$ $\mathbf{0.674}$ $\mathbf{0.586}$ $\mathbf{0.557}$ $\mathbf{0.616}$ $\mathbf{0.611}$
VI-E Results’ Discussion
We believe the strength of the proposed approach lies in the exploitation of the complex nature of SAR images for hindering the manipulation traces. Indeed, the speckle injection step contributes in widening the frequency spectrum of the manipulated image and conceals the subtle editing traces searched by forensic detectors. Though it is true that noise addition operations like those tested in Section VI-D are enlarging the spectrum as well, the speckle injection step is adding a complex phase term which is paramount to recreate the effect of a SAR acquisition. All the other global operations do not consider processing complex images, as they act only on the magnitude component.
@@ VII Conclusions
We evaluate the efficacy of the suggested counter-forensic method across a range of scenarios, analyzing different manipulation techniques. We experiment on data undergone to automatic and hand-made local manipulations, considering also the alarming scenario of military vehicles insertion. The results obtained demonstrate that our developed attack effectively eradicates any evidence of manipulation, even fooling sophisticated forensic detection systems. Moreover, the visual quality and the SAR properties of the attacked images are maintained, to the extent that even an expert SAR scientist would unlikely notice the difference with the original data.

### v2

@@ VI-D Realistic Attacks
Notice that the attacked spliced images still carry semantic traces of the inserted vehicle: in correspondence of the tank position, the scattered amplitude has a visible peak (see the fifth column). Despite this fact, the detectors are blind to the attacked image, and their heatmap cannot distinguish the splicing area from the surrounding pixels.
VI-E Results’ Discussion
@@
Figure 9: Examples of the proposed counter-forensic attack applied to manipulated amplitude images by exploiting military vehicles selected from MSTAR dataset. From left to right: the donor SAR amplitude image of the vehicle; the manipulated image in presence of attack; close-up of the pristine image of the area inside the white box; close-up of the manipulated image in absence of attack; close-up of the manipulated image in presence of attack; close-up of the TruForS detector heatmap in absence of attack; close-up of the TruForS detector heatmap in presence of attack.
@@ VII Conclusions
We evaluate the efficacy of the suggested counter-forensic method across a range of scenarios, analyzing different manipulation techniques. We experiment on data undergone to automatic and hand-made local manipulations, considering also the alarming scenario of military vehicles insertion. The results obtained demonstrate that our developed attack effectively eradicates any evidence of manipulation, even fooling sophisticated forensic detection systems. Moreover, the visual quality and the SAR properties of the attacked images are maintained. We also compare the proposed attack with another counter-forensic solution available in the literature. Our attack demonstrates better reconstruction properties in terms of SAR-related features and significantly stronger anti-forensic capabilities.
Possible Future Directions. In our proposed work, we limit the attack to concealing manipulation traces by injecting speckle noise from the same SAR system used for acquiring the original images. However, realistic scenarios could include situations where attackers might need to disguise a manipulated image as a pristine one captured by another SAR system. We believe that, in principle, our proposed methodology would still work, given that the other satellite chosen as the “target” system has very similar characteristics (acquisition mode, processing pipeline, etc.) with respect to the original one used for acquisitions. We reserve the possibility of exploring this in future studies.
Another interesting aspect to investigate is the transferability of the proposed attack to other detection models, such as SAR Automatic Target Recognition (ATR) models, which are commonly used to extract remote sensing information and have already revealed several vulnerabilities to adversarial attacks [44, 45]. Future works will focus on exploring whether our counter-forensic attack, designed for localization detectors, can also be effective in deceiving ATR models.
[…]